horse1234:

Sara borrowed $500 for four years at 3 percent interest, compounded annually. Use the formula to calculate the total amount she will have paid back.

Karla:

A=P(1+r/n)^nt P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year For your question A=? P=500 T=4 n=1 r=3% Put the values in equation A=500(1+3%/1)^4*1 A=500(1+3%)^4 A=500(1.03)^4 A=562.7